OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how to use the distributive property to simplify the expression below. 3(4 + x) Part 2: Write the simplified expression.

OpenStudy (mysesshou):

The 3 is multiplied by the contents in the parentheses. The 3 is multiplied to each term within the parentheses. This gives 3 times 4 and also 3 times x. 3*(4+x) =3*4 + 3*x =12 + 3x

OpenStudy (anonymous):

okai the word part is part one ritee? and the numberics is part two

OpenStudy (mysesshou):

yes, that would fit what they ask this part here 3*(4+x) =3*4 + 3*x =12 + 3x is the process to arrive at the simplified expression. the simplified expression is 12+3x

OpenStudy (mysesshou):

You may want to put the sentences into your own words so it shows that you understand it.